OpenStudy (anonymous):

\frac{ x^3y^5 }{ (x-6)(x+1) }times \frac{ (2x-3)(x-6) }{ x^6y^5 }

OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ x^2-5x-6 }\times \frac{ 2x^2-15x+18 }{ x^6y^5 }\] \frac{ x^3y^5 }{ (x-6)(x+1) }times \frac{ (2x-3)(x-6) }{ x^6y^5 }

OpenStudy (anonymous):

Simplify.